Kate has $13.65 in nickels, dimes, and quarters in her purse.
If the number of dimes is three more than the number of nickels, and the number of quarters is four times the number of dimes, how many dimes does Kate have in her purse?
Thank you in advanced. If you could please show let statements and work. :)

Answer by scott8148(6628)

let n=nickels , d=dimes , q=quarters

"Kate has $13.65 in nickels, dimes, and quarters" ___ 5n + 10d + 25q = 1365

"number of dimes is three more than the number of nickels" ___ d = n + 3 ___ d - 3 = n

"number of quarters is four times the number of dimes" ___ q = 4d

substituting ___ 5(d - 3) + 10d + 25(4d) = 1365

5d - 15 + 10d + 100d = 1365 ___ 115d = 1380